fx=3x squared -18x+4 find the vector, the line of symmetry and max. I have tried every thing from foil to factoring.
Answer by nerdybill(7384)


.
Everything you need to know about vertex form can be found at:
http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php
.
f(x)=3x^2 -18x+4
Group the x terms:
f(x) = (3x^2 -18x) + 4
Factor out a 3:
f(x) = 3(x^2 -6x) + 4
Complete the square:
f(x) = 3(x^2 -6x + ___ ) + 4
f(x) = 3(x^2 -6x + 9 ) + 4 - 18
f(x) = 3(x+3)^2 - 14
.
Vertex: (-3,-14)
Axis of symmetry: x= (-b)/(2a) = (18)/(2*3) = 3
x = 3
Max: Since the coefficient associated with the x^2 term is POSITIVE -- you will have a MINIMUM not a maximum. Minimum is at -14.